# Discrete Solution Operator Learning for Geometry-Dependent PDEs

### v1

@@ 2.3 Geometry-dependent linear elasticity problem
Fig. 4: Discrete solution operator learning (DiSOL) results for geometry-dependent 2D linear elasticity. a, Problem formulation and data representation for the linear elastic solid mechanics problem. The input to the model consists of five discrete channels defined on a fixed Cartesian grid, encoding the geometry mask, displacement boundary conditions in the horizontal and vertical directions, and externally applied force components. The model outputs the corresponding two-component displacement field $(\hat{u}_{x},\hat{u}_{y})$ defined on the same grid. b, Representative ID test results. Ground-truth displacement fields are compared with predictions from the proposed discrete solution operator (DiSOL) for two representative geometries and loading configurations. DiSOL accurately reconstructs both horizontal and vertical displacement components, capturing the global deformation patterns and boundary-induced responses without introducing spurious oscillations or rigid-body artifacts. c, OOD generalization results involving unseen geometries with increased complexity, including internal holes, topological changes, and altered boundary condition configurations. Despite the substantial changes in geometric structure and boundary constraints, DiSOL preserves the overall deformation modes and displacement distributions, maintaining consistency with the reference solutions for both displacement components. See Supplementary Information A.3, B and E.3 for full setup and extended statistics.
@@
Figure 4b presents representative ID test results (more results are presented in Supplementary Information E.3). DiSOL accurately reconstructs the spatial distributions of both displacement components, with predicted patterns closely matching the reference results in terms of global deformation patterns, boundary-induced displacement responses, and local gradient variations. Notably, no common non-physical artifacts, such as spurious oscillations, rigid-body drift, or inconsistencies between displacement components, are observed. This indicates that, within the training distribution, DiSOL successfully captures the discrete solution structure jointly determined by geometry, boundary conditions, and external loading.
@@ 2.4 Geometry-dependent thermal conduction problem
Fig. 5: Discrete solution operator learning (DiSOL) for the heat conduction problem. a, Problem formulation and data representation. The model input consists of discrete channels defined on a fixed Cartesian grid, encoding the geometry mask, boundary condition locations, source term distribution, and initial condition (IC). The time coordinate is provided as an additional input channel. The model predicts the temperature field at queried times $t$; during training, $t$ is sampled from $0,20]\,\mathrm{s}$. b, Representative ID test results at $t=2\,\mathrm{s}$, $10\,\mathrm{s}$, and $20\,\mathrm{s}$, showing close agreement between ground-truth solutions and DiSOL predictions. c, Zero-shot OOD evaluation combining geometric extrapolation and temporal forecasting. The model is tested on an unseen geometry with internal voids and altered boundary configurations. Predictions at $t=30\,\mathrm{s}$, $40\,\mathrm{s}$, and $50\,\mathrm{s}$ extend beyond the training temporal interval, corresponding to genuine future forecasting scenarios. Color bars indicate normalized temperature magnitude. See Supplementary Information A.4, B and E.4 for full setup and extended statistics.
@@
As in the previous cases, discretized geometry, boundary conditions, source terms, and the initial condition are supplied as discrete input channels to the model (Fig. 5a). To account for temporal evolution, the time coordinate is incorporated as an additional input channel and provided in parallel with the spatial inputs. Under this formulation, the learning task corresponds to approximating a discrete spatiotemporal solution operator defined on geometry-dependent computational domains. The model is trained on trajectories spanning the temporal interval $t\in[0,20]\,\mathrm{s}$.
@@
We further assess generalization under a zero-shot OOD setting (Figs. 5c and S15), where multiple unseen factors are introduced simultaneously. These include increased geometric complexity with internal voids, altered and discontinuous boundary configurations, and higher-frequency source and initial condition fields. In addition to spatial extrapolation, this test also probes temporal extrapolation: predictions at $t=30\,\mathrm{s}$, $40\,\mathrm{s}$, and $50\,\mathrm{s}$ extend beyond the training temporal interval and therefore constitute genuine future forecasting rather than interpolation. Despite this combined OOD setting in geometry, boundary conditions, source and initial fields, and time, DiSOL remains stable and produces solution patterns that closely match the numerical reference solutions across all evaluated time instances.
@@ 3 Discussion
In this work, we introduced a discrete solution operator learning (DiSOL) framework for geometry-dependent partial differential equations. Rather than approximating a global mapping between continuous function spaces, DiSOL operates directly on discretized computational domains and learns to assemble local solution contributions conditioned on geometry and boundary information. Through a series of benchmark problems—including scalar elliptic equations, transport-dominated advection–diffusion, vector-valued linear elasticity, and time-dependent heat conduction—we demonstrated that this discrete operator formulation enables robust generalization across geometry-dependent problem settings.
A central insight emerging from our results is that the primary challenge in geometry-dependent PDE learning does not stem from model capacity or optimization, but from the nature of geometric variation itself. Importantly, this behavior should not be interpreted as a fundamental limitation of continuous neural operator learning. Specifically, in many practical settings, changes in geometry induce discrete and non-smooth structural variations—such as alterations in domain topology, boundary classification, or active computational regions—that fundamentally modify how local contributions are assembled. When such geometry-induced effects dominate problem complexity, the mapping from problem specification to solution patterns is no longer well described as interpolation within a smooth function-space operator. Continuous operator methods remain highly effective when variations in problem specification, including geometry, can be treated as smooth perturbations of an underlying function-space mapping.
@@
This viewpoint provides a unified explanation for the consistent performance of DiSOL observed across all studied cases, ranging from scalar elliptic problems to transport-dominated dynamics, vector-valued elasticity, and spatiotemporal evolution. In each setting, performance degradation of continuous operator baselines is closely associated with geometry-induced discrete structural changes, rather than with increased physical complexity or stronger coupling alone. From a broader perspective, discrete and continuous operator learning paradigms should be viewed as complementary rather than competing: continuous operators offer elegant representations for problems dominated by smooth parametric variation, whereas discrete operator formulations are more naturally suited to cases in which geometry plays a central and structurally discrete role. The present study is limited to fixed-resolution discrete representations, and extending the framework to variable-resolution discretizations, incorporating stricter physical constraints, or developing hybrid formulations that combine discrete and continuous operator principles represent promising directions for future work.
@@ 4.1 Notation for normalized solutions
In this work, we let $U_{h}$ denote the ground-truth solution field with physical amplitude on the embedded grid (scalar for Poisson/advection-diffusion/heat conduction problems and vector-valued for elasticity). We define a per-sample amplitude $u_{\lim}$ (Supplementary Information. D, e.g., $u_{\lim}=\max_{x}|U_{h}(x)|$) and the dimensionless solution pattern $u_{h}:=U_{h}/u_{\lim}$, so that $\max_{x}|u_{h}(x)|=1$. All models (DiSOL and all baselines) are trained to predict the normalized pattern $\hat{u}_{h}$. If an absolute-amplitude prediction is required, it can be reconstructed as $\hat{U}_{h}=\hat{u}_{\lim}\,\hat{u}_{h}$ using an optional amplitude regressor (Supplementary Sec. D).
@@ 4.2 Problem setup and discrete learning objective
We formulate geometry-dependent PDE solving as learning a mapping from a discretized problem specification to its (discretized) solution on a fixed Cartesian grid. Each problem instance is represented by a multi-channel tensor $\boldsymbol{X}\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$ that encodes (i) a binary geometry mask $m\in\{0,1\}^{H\times W}$ indicating the computational domain $\Omega_{h}$, (ii) boundary-condition indicators on the embedded grid, and (iii) problem-dependent driving terms (e.g., sources, loads, or initial conditions). The goal is to learn an operator $\mathcal{F}_{\theta}$ that predicts the corresponding normalized solution pattern (or vector-valued patterns for elasticity) on the same grid
@@ 5 Acknowledgements
Support from the National Natural Science Foundation of China (Grant nos. 12032014 and T2488101) is acknowledged (X.-Q. Feng). Support from the National Natural Science Foundation of China (Grant no. 12502234) is acknowledged (J. Bai).

### v2

@@ 2.3 Geometry-dependent linear elasticity problem
Fig. 4: Discrete solution operator learning (DiSOL) results for geometry-dependent 2D linear elasticity. a, Problem formulation and data representation for the linear elastic solid mechanics problem. The input to the model consists of five discrete channels defined on a fixed Cartesian grid, encoding the geometry mask, displacement boundary conditions in the horizontal and vertical directions, and externally applied force components. The model outputs the corresponding two-component displacement field $(\hat{u}_{x},\hat{u}_{y})$ defined on the same grid. b, Representative ID validation results. Ground-truth displacement fields are compared with predictions from the proposed discrete solution operator (DiSOL) for two representative geometries and loading configurations. DiSOL accurately reconstructs both horizontal and vertical displacement components, capturing the global deformation patterns and boundary-induced responses without introducing spurious oscillations or rigid-body artifacts. c, OOD generalization results involving unseen geometries with increased complexity, including internal holes, topological changes, and altered boundary condition configurations. Despite the substantial changes in geometric structure and boundary constraints, DiSOL preserves the overall deformation modes and displacement distributions, maintaining consistency with the reference solutions for both displacement components. See Supplementary Information A.3, B and E.3 for full setup and extended statistics.
@@
Figure 4b presents representative ID validation results (more results are presented in Supplementary Information E.3). DiSOL accurately reconstructs the spatial distributions of both displacement components, with predicted patterns closely matching the reference results in terms of global deformation patterns, boundary-induced displacement responses, and local gradient variations. Notably, no common non-physical artifacts, such as unphysical oscillations, rigid-body drift, or inconsistencies between displacement components, are observed. This indicates that, within the training distribution, DiSOL successfully captures the discrete solution structure jointly determined by geometry, boundary conditions, and external loading.
@@ 2.4 Geometry-dependent thermal conduction problem
Fig. 5: Discrete solution operator learning (DiSOL) for the heat conduction problem. a, Problem formulation and data representation. The model input consists of discrete channels defined on a fixed Cartesian grid, encoding the geometry mask, boundary condition locations, source term distribution, and IC. The time coordinate is provided as an additional input channel. The model predicts the temperature field at queried times $t$; during training, $t$ is sampled from $0,20]\,\mathrm{s}$. b, Representative ID validation results at $t=2\,\mathrm{s}$, $10\,\mathrm{s}$, and $20\,\mathrm{s}$, showing close agreement between ground-truth solutions and DiSOL predictions. c, Zero-shot OOD evaluation combining geometric extrapolation and temporal forecasting. The model is tested on an unseen geometry with internal voids and altered boundary configurations. Predictions at $t=30\,\mathrm{s}$, $40\,\mathrm{s}$, and $50\,\mathrm{s}$ extend beyond the training temporal interval, corresponding to genuine future forecasting scenarios. Color bars indicate normalized temperature magnitude. See Supplementary Information A.4, B and E.4 for full setup and extended statistics.
@@
As in the previous cases, discretized geometry, boundary conditions, source terms, and the initial condition (IC) are supplied as discrete input channels to the model (Fig. 5a). To account for temporal evolution, the time coordinate is incorporated as an additional input channel and provided in parallel with the spatial inputs. Under this formulation, the learning task corresponds to approximating a discrete spatiotemporal solution operator defined on geometry-dependent computational domains. The model is trained on trajectories spanning the temporal interval $t\in[0,20]\,\mathrm{s}$.
@@
We further assess generalization under a zero-shot OOD setting (Figs. 5c and S15), where multiple unseen factors are introduced simultaneously. These include increased geometric complexity with internal voids, altered and discontinuous boundary configurations, and higher-frequency sources and IC fields. In addition to spatial extrapolation, this test also probes temporal extrapolation: predictions at $t=30\,\mathrm{s}$, $40\,\mathrm{s}$, and $50\,\mathrm{s}$ extend beyond the training temporal interval and therefore constitute genuine future forecasting rather than interpolation. Despite this combined OOD setting in geometry, boundary conditions, sources and IC fields, and time, DiSOL remains stable and produces solution patterns that closely match the numerical reference solutions across all evaluated time instances.
@@ 3 Discussion
In this work, we introduced the DiSOL framework for geometry-dependent partial differential equations. Rather than approximating a global mapping between continuous function spaces, DiSOL operates directly on discretized computational domains and learns to assemble local solution contributions conditioned on geometry and boundary information. Through a series of benchmark problems—including the Poisson equation, transport-dominated advection-diffusion, vector-valued linear elasticity, and time-dependent heat conduction—we demonstrated that this discrete operator formulation enables robust generalization across geometry-dependent problem settings.
A central insight emerging from our results is that the primary challenge in geometry-dependent PDE learning does not stem from model capacity or optimization, but from the nature of geometric variation itself. Importantly, this behavior should not be interpreted as a fundamental limitation of continuous neural operator learning. Specifically, in many practical settings, changes in geometry induce discrete and non-smooth structural variations—such as alterations in domain topology, boundary classification, or active computational regions—that fundamentally modify how local contributions are assembled. When such geometry-induced effects dominate problem complexity, the mapping from problem specification to solution field is no longer well described as interpolation within a smooth function-space operator. Continuous operator methods remain highly effective when variations in problem specification, including geometry, can be treated as smooth perturbations of an underlying function-space mapping.
@@
DiSOL’s advantage can be understood through a discrete-operator perspective. The local contribution operator learns reusable, stencil-like transformation rules that map neighborhood-level specifications (geometry, boundary activation, and sources) to local responses, analogous to local discretization stencils shared across the domain. The multiscale assembly pathway then aggregates these local contributions across resolutions, resembling an implicit form of domain decomposition in which coarse features convey long-range coupling while fine features refine boundary-adjacent structure. Under geometry shifts, the procedure (local rules + assembly) remains invariant, while only the set of activated locations changes; this separation helps preserve local validity and enables robust reassembly under new global connectivity and boundary configurations. A more detailed conceptual analysis is provided in Supplementary Information G.
This viewpoint provides a unified explanation for the consistent performance of DiSOL observed across all studied cases, ranging from scalar Poisson equation to transport-dominated dynamics, vector-valued elasticity, and spatiotemporal evolution. In each setting, performance degradation of continuous operator baselines is closely associated with geometry-induced discrete structural changes, rather than with increased physical complexity or stronger coupling alone. From a broader perspective, discrete and continuous operator learning paradigms should be viewed as complementary rather than competing: continuous operators offer elegant representations for problems dominated by smooth parametric variation, whereas discrete operator formulations are more naturally suited to cases in which geometry plays a central and structurally discrete role. The present study is limited to fixed-resolution discrete representations; extending the framework to variable-resolution discretizations (with additional cross-resolution results reported in Supplementary Information E.5), incorporating stricter physical constraints, or developing hybrid formulations that combine discrete and continuous operator principles are promising directions for future work. We also document challenging elasticity regimes where DiSOL exhibits localized artifacts (Supplementary Information F), indicating that stronger boundary-consistent constraints and improved representations may be required in these cases.
@@ 4.1 Notation for normalized solutions
In this work, we let $U_{h}$ denote the ground-truth solution field with physical amplitude on the embedded grid (scalar for Poisson/advection-diffusion/heat conduction problems and vector-valued for elasticity). We define a per-sample amplitude $u_{\lim}$ (Supplementary Information D, e.g., $u_{\lim}=\max_{x}|U_{h}(x)|$) and the dimensionless solution pattern $u_{h}:=U_{h}/u_{\lim}$, so that $\max_{x}|u_{h}(x)|=1$. All models (DiSOL and all baselines) are trained to predict the normalized pattern $\hat{u}_{h}$. If an absolute-amplitude prediction is required, it can be reconstructed as $\hat{U}_{h}=\hat{u}_{\lim}\,\hat{u}_{h}$ using an optional amplitude regressor (Supplementary Sec. D).
@@ 4.2 Problem setup and discrete learning objective
We formulate geometry-dependent PDE solving as learning a mapping from a discretized problem specification to its (discretized) solution on a fixed Cartesian grid. Each problem instance is represented by a multi-channel tensor $\boldsymbol{X}\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$ that encodes (i) a binary geometry mask $m\in\{0,1\}^{H\times W}$ indicating the computational domain $\Omega_{h}$, (ii) boundary-condition indicators on the embedded grid, and (iii) problem-dependent driving terms (e.g., sources, loads, or IC fields). The goal is to learn an operator $\mathcal{F}_{\theta}$ that predicts the corresponding normalized solution pattern (or vector-valued patterns for elasticity) on the same grid
@@ 5 Acknowledgements
Support from the National Natural Science Foundation of China (Grant nos. T2488101 and 12032014) is acknowledged (X.-Q. Feng and J. Bai). Support from the National Natural Science Foundation of China (Grant no. 12502234) is acknowledged (J. Bai).